OpenStudy (anonymous):

Can someone please help me finish this pre-calc question??? Express the complex number in trigonometric form. 2 - 2i This is what I have so far: r=absolute value(2-2i) r= sqrt((2)^2+(-2)^2) r=sqrt(8) r=2sqrt(2) theta(ref)=tan^-1*absolute value(2/-2) theta(ref)=pi/2 I dont know how to get the value of theta alone to complete this problem. What do i do next? Please help!

OpenStudy (anonymous):

\[2-2i\]\[r=\sqrt{2^2+(-2)^2}\]\[\tan^{-1} = (\frac{ -2 }{ 2 } ) = ?\]

OpenStudy (anonymous):

you arent suppose to use radian mode for polar form

OpenStudy (anonymous):

thats why you got pi/2

OpenStudy (anonymous):

I thought \[\tan^{-1} =\frac{ a }{ b }\]

OpenStudy (anonymous):

no its b/a

OpenStudy (anonymous):



OpenStudy (anonymous):

better if i draw it out?

OpenStudy (anonymous):

Yes. After switching it to b/a i got theta(ref)=-1

OpenStudy (anonymous):

find out whats tan^1(-1)

OpenStudy (anonymous):

tan^1(-1) or tan^-1(1)? tan^1(-1) is -1.55 or -tan(1) http://www.wolframalpha.com/input/?i=tan%5E1%28-1%29

OpenStudy (anonymous):

\[\tan^{-1} (-1)\]

OpenStudy (anonymous):

Okay, yeah tan^-1(-1) is -pi/4 or -45 degrees

OpenStudy (anonymous):

now put that in 360 degrees form

OpenStudy (anonymous):

find the angle that is coterminal to -45 degrees? Thats 315, right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

you know how to make the polar form from there right?

OpenStudy (anonymous):

oh wait its trigo form

OpenStudy (anonymous):

Yea. I was stuck at theta(ref)= -1. I dont know what to do next

OpenStudy (anonymous):

ok great you have the angles now you can form whatever it is

OpenStudy (anonymous):

you should get \(\theta\) instantly from your eyeballs